OpenStudy (anonymous):

Find the domain and range of a real function f(x) =1/(1-x^2)?

OpenStudy (anonymous):

domain = R-[-1 , 1] is that correct

OpenStudy (anonymous):

then wat is the range??

OpenStudy (anonymous):

hw to find it

OpenStudy (helder_edwin):

solve for x the equation \[ y=f(x)=\frac{1}{1-x^2} \] you'll get something like x=g(y) and do for this whatever you did to find the domain of f

OpenStudy (experimentx):

\[ 1 \over 1 - x^2\]

OpenStudy (experimentx):

what values of x is not possible? and what outcome is not possible?

OpenStudy (anonymous):

0 is not possible

OpenStudy (anonymous):